OpenStudy (anonymous):

the sum of two consecutive odd integers is -148 find the two integers

OpenStudy (gw2011):

Let x=one of the odd integers Then x-2=the other odd integer x+(x-2)=-148 x+x-2=-1482 2x-2=-148 2x=-148+2 2x=-146 x=-73 x-2=-75 Therefore, the two consecutive odd integers are -73 and -75